undeadknight26 (undeadknight26):

Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1).

undeadknight26 (undeadknight26):

y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3) I think it is the first one.

OpenStudy (zale101):